# Strategic Learning and Trading in Broker-Mediated Markets

### v1

@@ 3.4   The optimal adaptive strategy
The following theorem solves the adaptive control problem of the informed trader.
@@
The HJB equation (3.10) admits a unique solution in $C^{1,2}([0,T],\mathbb{R}^{|\mathfrak{x}|})$ in the form of
@@
where $g_{i}^{I}$, $i=0,1,2$ are defined in Lemma 1. Moreover, there exists a constant $c<\infty$ such that
@@
$\eta^{*}_{t}=\frac{z_{1}^{I}(t)\,\alpha_{t}+z_{2}^{I}(t)\,\hat{\nu}_{t}+2\,g_{2}^{I}(t)\,Q^{I*}_{t}}{2\,\mathfrak{b}}\,.$ (3.26)
where $z_{1}^{I},z_{2}^{I}\geq 0$ and $g_{2}^{I}<0$.
Proof.
Observe that (3.24) satisfies the HJB equation in (3.10). Moreover, by substituting $g_{i}^{I}$ for $i=0,1,2$ from Lemma 1, the solution to the HJB equation is a quadratic polynomial in terms of $\mathfrak{x}$ with time-dependent $C^{1}([0,T])$ coefficients; thus, the solution is bounded by a quadratic growth. By Theorem 3.5.2 (Verification Theorem) in Pham, (2009), the trader’s strategy (3.26) is indeed the optimal Markovian control. The last part of the theorem is as follows: $g_{2}^{I}$ is negative from Lemma 1.
Next, by solving for the ODE (3.21), we obtain
$z_{2}^{I}(t)=\int_{t}^{T}\mathfrak{p}\,\exp\left\{\int_{t}^{s}\frac{g_{2}^{I}(u)}{2\,\mathfrak{b}}-\theta^{B}\,\mathrm{d}u\right\}\,\mathrm{d}s\geq 0\quad\forall t\in[0,T]\,,$ (3.27)
which shows that $z_{2}^{I}$ is non-negative. The same argument follows for $z_{1}^{I}$ by solving (3.20). Observe that the equation above is linear in $\mathfrak{p}$. ∎
@@
If the informed trader were to observe $\nu$ and if $\nu$ was indeed an OU process, then the HJB equation of such full-information control problem would differ from (3.10) in the coefficients of $\partial_{\hat{\nu},\hat{\nu}}H^{I}$, $\partial_{s,\hat{\nu}}H^{I}$, and $\partial_{\alpha,\hat{\nu}}H^{I}$ only. From here, we use a version of Lemma 1 to show that the equivalent coefficients to $g_{2}^{I}$ and $g_{1}^{I}$ we obtain in such full-information case are exactly the same but changing $\hat{\nu}$ for $\nu$. This is the separation principle discussed in Section 9.3 of Bensoussan, (2018).
@@ 4.1   Learning equations
However, the broker does not observe $\alpha$, nor can she construct the filter $\hat{\nu}_{t}$ because the innovation process requires knowledge of $\alpha$. Thus, the broker replaces the signal $\alpha$ in (4.1) with an estimate $\hat{\alpha}$. Below, Section 4.1.1 assumes that the broker employs midprices to construct this estimate and in Section 4.1.2 the broker employs the trading speed of the informed trader as the source of information.
@@ 4.1.1 Learning from prices
Here, the broker assumes that her competitor’s trading speed takes the form
$\eta^{*}_{t}=\frac{z_{1}^{I}(t)\,\hat{\alpha}_{t}+z_{2}^{I}(t)\,\nu_{t}+2\,g_{2}^{I}(t)\,Q^{I*}_{t}}{2\,\mathfrak{b}}\,,$ (4.2)
where she replaces the estimate $\hat{\nu}$ with her true trading speed $\nu$ in (4.1)—this is equivalent to assuming that the filter $\hat{\nu}_{t}$ of the informed trader matches the true $\nu_{t}$. This latter assumption may be conservative, as perfect knowledge gives an obvious advantage to the informed trader. However, we show in the numerical examples of Section 5 that the broker can strategically influence (manipulate) the informed trader through this component, i.e., through knowing and exploiting that the informed trader’s strategy is based on that of the broker.
@@ 4.1.2 Learning from trading flow
The broker is counterparty to the trades of her clients; the informed and uninformed trader, for which she charges proportional transaction costs. Thus, she observes the trading flow of the informed trader. Here, we assume that the broker employs this trading flow as a source of information from which she obtains an estimate of the private signal $\alpha$. Here, $\eta^{*}$ is a measurable stochastic process that carries information about the unknown signal $\alpha$; c.f. with the assumptions in Section 4.1.1. The broker assumes the parametric form (4.1), so she observes the process ${\gamma}_{t}:=\eta^{*}_{t}-f_{3}(t)\,Q^{I*}_{t}=f_{1}(t)\,\alpha_{t}+f_{2}(t)\,\hat{\nu}_{t}$. We compute the dynamics of ${\gamma}_{t}$ to obtain
$\begin{split}\mathrm{d}{\gamma}_{t}&=\quad\left[f_{1}^{\prime}(t)\,\alpha_{t}-\kappa^{\alpha}\,f_{1}(t)\,\alpha_{t}+f_{2}^{\prime}(t)\,\hat{\nu}_{t}-\theta^{B}\,f_{2}(t)\,\hat{\nu}_{t}\right]\,\mathrm{d}t\\ &\qquad\qquad+\left[\sigma^{\alpha}\,f_{1}(t)\,\mathrm{d}W_{t}^{\alpha}+\frac{f_{2}(t)}{(\sigma^{S})^{2}}\,\mathfrak{p}\,\mathbb{V}^{I}_{t}\,\left(\mathrm{d}Y_{t}-\mathfrak{p}\,\hat{\nu}_{t}\,\mathrm{d}t\right)\right]\\ &=\quad\Bigg{[}\Big{(}f_{1}^{\prime}(t)-\kappa^{\alpha}\,f_{1}(t)\Big{)}\,\alpha_{t}+\left(f_{2}^{\prime}(t)-\theta^{B}\,f_{2}(t)-\frac{\mathfrak{p}^{2}\,\mathbb{V}^{I}_{t}\,f_{2}(t)}{\left(\sigma^{S}\right)^{2}}\right)\,\hat{\nu}_{t}+\frac{\mathfrak{p}^{2}\,\mathbb{V}^{I}_{t}\,f_{2}(t)}{\left(\sigma^{S}\right)^{2}}\,\nu_{t}\Bigg{]}\,\mathrm{d}t\\ &\qquad\qquad+\left[\sigma^{\alpha}\,f_{1}(t)\,\mathrm{d}W_{t}^{\alpha}+\frac{\mathfrak{p}\,\mathbb{V}^{I}_{t}\,f_{2}(t)}{\sigma^{S}}\,\mathrm{d}W_{t}^{S}\right].\end{split}$ (4.7)
The broker knows that
so she substitutes this in the above equation to obtain
$\begin{split}\mathrm{d}{\gamma}_{t}&=\left[\mathfrak{G}^{\alpha}(t)\,\alpha_{t}+\mathfrak{G}_{0}(t)\,{\gamma}_{t}+\mathfrak{G}_{1}(t)\,\nu_{t}+\mathfrak{G}_{2}(t)\right]\,\mathrm{d}t\\ &\qquad\qquad\qquad\qquad+\left[\mathfrak{G}_{3}(t)\,\mathrm{d}W_{t}^{\alpha}+\mathfrak{G}_{4}(t)\,\mathrm{d}W_{t}^{S}\right]\,,\end{split}$ (4.8)
where, using (3.20)–(3.21), one obtains
$\mathfrak{G}_{0}(t)$ $=-\theta^{B}-\frac{\mathfrak{p}^{2}\,\mathbb{V}^{I}_{t}}{\left(\sigma^{S}\right)^{2}}+\frac{f_{2}^{\prime}(t)}{f_{2}(t)}$ (4.9)
$=-\frac{\mathfrak{p}^{2}\,\mathbb{V}^{I}_{t}}{\left(\sigma^{S}\right)^{2}}-\frac{\mathfrak{p}}{2\,f_{2}(t)\,\mathfrak{b}}-\frac{f_{3}(t)}{2}\,,$ (4.10)
$\mathfrak{G}^{\alpha}(t)$ $=f_{1}^{\prime}(t)-f_{1}(t)\,\left(\kappa^{\alpha}+\mathfrak{G}_{0}(t)\right)$ (4.11)
$=-\frac{1}{2\,\mathfrak{b}}-f_{1}(t)\,\left(\frac{f_{3}(t)}{2}+\mathfrak{G}_{0}(t)\right)\,,$ (4.12)
$\mathfrak{G}_{1}(t)$ $=\frac{\mathfrak{p}^{2}\,\mathbb{V}^{I}_{t}\,f_{2}(t)}{\left(\sigma^{S}\right)^{2}},\quad\mathfrak{G}_{2}(t)=0\,,$ (4.13)
$\mathfrak{G}_{3}(t)$ $=\sigma^{\alpha}\,f_{1}(t),\quad\quad\mathfrak{G}_{4}(t)=\frac{\mathfrak{p}\,\mathbb{V}^{I}_{t}\,f_{2}(t)}{\sigma^{S}}\,.$ (4.14)
We define the process $\tilde{Z}_{t}:=\gamma_{t}/\mathfrak{G}_{5}(t)$ with $\mathfrak{G}_{5}(t):=\sqrt{\mathfrak{G}_{3}(t)^{2}+\mathfrak{G}_{4}(t)^{2}+2\,\rho\,\mathfrak{G}_{3}(t)\,\mathfrak{G}_{4}(t)}$ that follows the dynamics
$\begin{split}\mathrm{d}\tilde{Z}_{t}&=\quad\left[\frac{-\tilde{Z}_{t}\,\mathfrak{G}_{5}^{{}^{\prime}}(t)}{\mathfrak{G}_{5}(t)}+\frac{\mathfrak{G}^{\alpha}(t)}{\mathfrak{G}_{5}(t)}\,\alpha_{t}+\frac{\mathfrak{G}_{0}(t)}{\mathfrak{G}_{5}(t)}\,\gamma_{t}+\frac{\mathfrak{G}_{1}(t)}{\mathfrak{G}_{5}(t)}\,\nu_{t}+\frac{\mathfrak{G}_{2}(t)}{\mathfrak{G}_{5}(t)}\right]\,\mathrm{d}t\\ &\qquad\qquad+\frac{1}{\mathfrak{G}_{5}(t)}\,\left[\mathfrak{G}_{3}(t)\,\mathrm{d}W_{t}^{\alpha}+\mathfrak{G}_{4}(t)\,\mathrm{d}W_{t}^{S}\right]\,.\end{split}$ (4.15)
The coefficients in the above expression are well-defined when $t\to T$. Observe that
$\begin{split}\lim_{t\uparrow T}\tilde{Z}_{t}&=\lim_{t\uparrow T}\frac{f_{1}(t)}{\mathfrak{G}_{5}(t)}\,\alpha_{t}+\frac{f_{2}(t)}{\mathfrak{G}_{5}(t)}\,\hat{\nu}_{t}\\ &=\left(\lim_{t\uparrow T}\frac{f_{1}(t)}{\mathfrak{G}_{5}(t)}\right)\,\alpha_{T}+\left(\lim_{t\uparrow T}\frac{f_{2}(t)}{\mathfrak{G}_{5}(t)}\right)\,\hat{\nu}_{T}\,,\end{split}$ (4.16)
and from L’Hopital’s rule and the differential equations (3.20)–(3.21), it follows that $\lim_{t\uparrow T}{f_{1}(t)}/{\mathfrak{G}_{5}(t)}$ and $\lim_{t\uparrow T}{f_{2}(t)}/{\mathfrak{G}_{5}(t)}$ exist and are finite. This implies that $\tilde{Z}$ is a continuous square-integrable process. similarly, one also shows that the limits of ${\mathfrak{G}^{\alpha}(t)}/{\mathfrak{G}_{5}(t)}$, ${\mathfrak{G}_{1}(t)}/{\mathfrak{G}_{5}(t)}$, and ${\mathfrak{G}_{2}(t)}/{\mathfrak{G}_{5}(t)}$ exist and are finite as $t\uparrow T$.
On the other hand, observe that ${\mathfrak{G}_{5}^{{}^{\prime}}(t)}/{\mathfrak{G}_{5}(t)}$ and ${\mathfrak{G}_{0}(t)}/{\mathfrak{G}_{5}(t)}$ blow up as $t\uparrow T$. However, because $(\tilde{Z}_{t})_{t\in{\mathfrak{T}}}$ and $\big{(}\int_{0}^{t}\alpha_{s}\,{\mathfrak{G}^{\alpha}(s)}/{\mathfrak{G}_{5}(s)}+{\mathfrak{G}_{1}(s)}/{\mathfrak{G}_{5}(s)}\,\nu_{s}+{\mathfrak{G}_{2}(s)}/{\mathfrak{G}_{5}(s)}\,\mathrm{d}s\big{)}_{t\in{\mathfrak{T}}}$ are linear combinations of square-integrable progressively measurable continuous processes, the process $\big{(}\int_{0}^{t}{-\tilde{Z}_{s}\,\mathfrak{G}_{5}^{{}^{\prime}}(s)}/{\mathfrak{G}_{5}(s)}+\gamma_{s}\,{\mathfrak{G}_{0}(s)}/{\mathfrak{G}_{5}(s)}\,\mathrm{d}s\big{)}_{t\in{\mathfrak{T}}}$ is also square-integrable and progressively measurable. Next we define the main process that we use in the filtering framework of this section. The broker observes the process $Z_{t}$ defined as
$\begin{split}\mathrm{d}Z_{t}:&=\mathrm{d}\tilde{Z}_{t}-\mathfrak{f}_{t}\,\mathrm{d}t\\ &=\mathfrak{G}_{7}(t)\,\alpha_{t}\,\mathrm{d}t+\mathrm{d}\tilde{W}_{t}\,,\end{split}$ (4.17)
$\mathrm{d}\,\tilde{W}_{t}$ $:=\frac{1}{\mathfrak{G}_{5}(t)}\,\left[\mathfrak{G}_{3}(t)\,\mathrm{d}W_{t}^{\alpha}+\mathfrak{G}_{4}(t)\,\mathrm{d}W_{t}^{S}\right]\,,$ (4.18)
$\mathfrak{G}_{6}(t)$ $=-\mathfrak{G}_{5}^{{}^{\prime}}(t)/\mathfrak{G}_{5}(t)\,,$ (4.19)
$\mathfrak{G}_{7}(t)$ $=\mathfrak{G}^{\alpha}(t)/\mathfrak{G}_{5}(t),\quad\mathfrak{G}_{8}(t)=\mathfrak{G}_{0}(t)/\mathfrak{G}_{5}(t)\,,$ (4.20)
$\mathfrak{G}_{9}(t)$ $=\mathfrak{G}_{1}(t)/\mathfrak{G}_{5}(t),\quad\mathfrak{G}_{10}(t)=\mathfrak{G}_{2}(t)/\mathfrak{G}_{5}(t)\,.$ (4.21)
Use the ODEs (3.20), (3.21), and (3.4), to rewrite $\mathfrak{G}_{6}(t)$ as
$\mathfrak{G}_{6}(t)=-\frac{\mathfrak{G}_{3}^{{}^{\prime}}(t)[\mathfrak{G}_{3}(t)+\rho\,\mathfrak{G}_{4}(t)]+\mathfrak{G}_{4}^{{}^{\prime}}(t)[\rho\,\mathfrak{G}_{3}(t)+\mathfrak{G}_{4}(t)]}{\mathfrak{G}_{5}(t)^{2}}\,,$ (4.22)
where
$\mathfrak{G}_{3}^{{}^{\prime}}(t)$ $=\sigma^{\alpha}\left[-\frac{1}{2\,\mathfrak{b}}+\kappa^{\alpha}\,f_{1}(t)-\frac{f_{3}(t)\,f_{1}(t)}{2}\right]\,,$ (4.23)
$\mathfrak{G}_{4}^{{}^{\prime}}(t)$ $=\frac{\mathfrak{p}}{\sigma^{S}}\left[\mathbb{V}^{I}_{t}\,\left(-\frac{\mathfrak{p}}{2\,\mathfrak{b}}+\theta^{B}\,f_{2}(t)-\frac{f_{3}(t)\,f_{2}(t)}{2}\right)+f_{2}(t)\,\left((\sigma^{B})^{2}-2\,\theta^{B}\,\mathbb{V}^{I}_{t}-\frac{\mathfrak{p}^{2}\,(\mathbb{V}^{I}_{t})^{2}}{(\sigma^{S})^{2}}\right)\right]\,.$ (4.24)
The broker sets up the filtering framework
@@
where $Z$ is the observed process and $\alpha$ is the unobserved process. Finally, she takes $(\mathcal{Z}_{t})_{t\in{\mathfrak{T}}}$ as the augmented filtration generated by the process $Z$. The proposition below computes, explicitly, the dynamics of the new filter.
@@
where $I$ is an innovation process satisfying the equation
@@
and
The proof of Proposition 1, which is a slight extension of Corollary 3.38 in Bain and Crisan, (2009) for time-dependent coefficients, is in A.
4.2   The model
This section derives the closed-form optimal strategy of the broker when she employs the learning equations of Section 4.1.1, i.e., when she employs the midprice as the source of information. The derivation of the optimal adaptive strategy using the learning dynamics (4.28) of Section 4.1.2 is similar, and thus omitted. We remark that in the numerical experiments below, we study the performance of the broker’s strategy when she estimates the private signal according to (4.4)–(4.5) and also (4.27)–(4.28).
The broker assumes that the uninformed order flow $\xi$ is mean-reverting around zero and we write its dynamics as
@@ 4.2   The model
$\mathrm{d}X^{B,\nu}_{t}$ =$=\left\{-\nu_{t}\,\left(S_{t}+\mathfrak{a}\,\nu_{t}\right)+\eta^{*}_{t}\,\left(S_{t}+\mathfrak{b}\,\eta^{*}_{t}\right)+\xi_{t}\,\left(S_{t}+\mathfrak{c}\,\xi_{t}\right)\right\}\,\mathrm{d}t\,,\quad X^{B,\nu}_{0}=0\,,$ (4.33)
@@
The broker considers the following set of admissible controls for her trading strategy in the lit market:
[…]
For a given strategy $\nu\in{\mathcal{A}}^{B}$, she maximises the performance criterion
@@
where $\mathfrak{y}=(s,x^{B},q^{B},\hat{\alpha},\xi,q^{I})$, and the interpretation is similar to that of the informed trader. Thus, the control problem that the broker solves is
@@
4.3   Hamilton–Jacobi–Bellman equation
The HJB equation associated with the value function $H$ in (4.36) is
[…]
$-\kappa^{\alpha}\,\hat{\alpha}\,\partial_{\hat{\alpha}}H+\frac{1}{2}\,\left(\frac{\mathbb{V}^{B}_{t}+\rho\,\sigma^{S}\,\sigma^{\alpha}}{\sigma^{S}}\right)^{2}\,\partial_{\hat{\alpha},\hat{\alpha}}H+\left(\mathbb{V}^{B}_{t}+\rho\,\sigma^{S}\,\sigma^{\alpha}\right)\,\partial_{\hat{\alpha},s}H$ (4.38)
$-\nu\,\left(s+\mathfrak{a}\,\nu\right)\,\partial_{x^{B}}H+\eta^{*}\,\left(s+\mathfrak{b}\,\eta^{*}\right)\,\partial_{x^{B}}H+\xi\,\left(s+\mathfrak{c}\,\xi\right)\,\partial_{x^{B}}H$ (4.39)
$+\left(\nu-\eta^{*}-\xi\right)\,\partial_{q^{B}}H$ (4.40)
$+\left(\eta^{*}\right)\,\partial_{q^{I}}H-\kappa^{U}\,\xi\,\partial_{\xi}H+\frac{1}{2}\,\left(\sigma^{U}\right)^{2}\,\partial_{\xi\xi}H$ (4.41)
$-\left(\rho_{0}^{B}+\rho_{1}^{B}\,\mathbb{V}^{B}_{t}\right)\left(q^{B}\right)^{2}\bigg{\}}\,,$ (4.42)
with terminal condition

### v2

@@ 3.3  The optimal adaptive strategy
The above control problem has a linear-quadratic structure, hence, we are able to characterise the solution with a Riccati differential equation and linear ODEs. We show existence and uniqueness of these differential equations and use them to solve the control problem of the informed trader. The following theorem, whose proof is in the appendix, shows the solution to the adaptive control problem of the informed trader.
@@
The HJB equation (A.2) admits a unique solution in $C^{1,2}([0,T],\mathbb{R}^{|\mathfrak{x}|})$ in the form of
@@
where $g_{i}^{I}$, $i=0,1,2$ are defined in the appendix (Lemma 1). Moreover, there exists a constant $c<\infty$ such that
@@
$\eta^{*}_{t}=\frac{1}{2\,\mathfrak{b}}\Big({z_{1}^{I}(t)\,\alpha_{t}+z_{2}^{I}(t)\,\hat{\nu}_{t}+2\,g_{2}^{I}(t)\,Q^{I*}_{t}}\Big)\,.$ (3.11)
where $z^{I}_{1},z^{I}_{2}$ are deterministic functions defined in the appendix and they satisfy that $z_{1}^{I},z_{2}^{I}\geq 0$ and $g_{2}^{I}<0$.
@@
If the informed trader were to observe $\nu$ and if $\nu$ was indeed an OU process, then the HJB equation of such full-information control problem would differ from (A.2) in the coefficients of $\partial_{\hat{\nu},\hat{\nu}}H^{I}$, $\partial_{s,\hat{\nu}}H^{I}$, and $\partial_{\alpha,\hat{\nu}}H^{I}$ only. From here, we use a version of Lemma 1 to show that the equivalent coefficients to $g_{2}^{I}$ and $g_{1}^{I}$ we obtain in such full-information case are exactly the same but changing $\hat{\nu}$ for $\nu$. This is the separation principle discussed in Section 9.3 of Bensoussan (2018).
@@ 4.1  Learning equations
However, the broker does not observe $\alpha$, nor can she construct the filter $\hat{\nu}_{t}$ because the innovation process requires knowledge of $\alpha$. Thus, the broker replaces the signal $\alpha$ in (4.1) with an estimate $\hat{\alpha}$. Below, Section 4.1.1 assumes that the broker employs midprices to construct this estimate and in Section 4.1.2 the broker employs the trading speed of the informed trader as the source of information. One of our main results follows from comparing the performance of these two approaches.
@@ 4.1.1 Learning from prices
In this subsection the broker filters the signal from observing the price process. To do this, the broker replaces the unknown processes $\alpha$ and $\hat{\nu}$ (used in the informed trader’s strategy) with the known measurable processes $\hat{\alpha}$ and $\nu$, respectively. More precisely, the broker assumes that her competitor’s trading speed takes the form
$\eta^{*}_{t}=\frac{1}{2\,\mathfrak{b}}\Big({z_{1}^{I}(t)\,\hat{\alpha}_{t}+z_{2}^{I}(t)\,\nu_{t}+2\,g_{2}^{I}(t)\,Q^{I*}_{t}}\Big)=f_{1}(t)\,\hat{\alpha}_{t}+f_{2}(t)\,\nu_{t}+f_{3}(t)\,Q^{I*}_{t}\,.$ (4.2)
This latter assumption may be conservative, as perfect knowledge gives an obvious advantage to the informed trader. However, we show in the numerical examples of Section 5 that the broker can strategically influence (manipulate) the informed trader through this component, i.e., through knowing and exploiting that the informed trader’s strategy is based on that of the broker. The functions $f_{i}$ for $i=1,2,3$ are those induced by the second equality.
@@ 4.1.2 Learning from trading flow
In this subsection the broker employs the trading flow as a source of information from which she obtains an estimate of the private signal $\alpha$. Observe that $\eta^{*}$ is a measurable stochastic process that carries information about the unknown signal $\alpha$; c.f. with the assumptions in Section 4.1.1. Using the parametric form (4.1), it follows that the broker observes the process ${\gamma}_{t}:=\eta^{*}_{t}-f_{3}(t)\,Q^{I*}_{t}=f_{1}(t)\,\alpha_{t}+f_{2}(t)\,\hat{\nu}_{t}$. Observe that this implicitly assumes that the broker knows the value of the parameters in the dynamics of the alpha signal. Below, in the numerical section we carry out robustness checks to test how the performance of the broker’s strategy changes as the broker mispecifies the values of these model parameters.
Using $\gamma_{t}$ we define a scaled version $\tilde{Z}^{\textup{alt}}_{t}$ and a shifted version $Z^{\textup{alt}}_{t}$. For the sake of presentation we report the details in the appendix. The key result is that one is left with the filtering framework
@@
where $Z^{\textup{alt}}$ is the observed process, $\mathfrak{G}_{7}$ is a deterministic function reported in B.1, $\tilde{W}$ is a $(\mathbb{P},(\mathcal{F}_{t})_{t\in{\mathfrak{T}}}$-Brownian motion, and $\alpha$ is the unobserved process. Finally, she takes $(\mathcal{Z}_{t})_{t\in{\mathfrak{T}}}$ as the augmented filtration generated by the process $Z^{\textup{alt}}$. The proposition below computes, explicitly, the dynamics of the new filter.
@@
where $\mathfrak{K}(t)$ is a deterministic function and $I$ is an innovation process satisfying the equation
@@
which is a $(\mathbb{P},(\mathcal{Z}_{t})_{t\in{\mathfrak{T}}})$-Brownian motion.
The proof of Proposition 1 is in B.2. As the noise flow $\xi$ is independent of either $W^{S}$ and $W^{\alpha}$, the above result would not change if we replace the filtration $(\mathcal{Z}_{t})_{t\in{\mathfrak{T}}}$ with the filtration ${\mathcal{F}}^{B,{\textup{alt}}}_{t}:=\sigma(Z^{\textup{alt}}_{s},\xi_{s}:\,s\in[0,t])\vee{\mathcal{N}}$.
4.2  The control problems of the broker
This section derives the closed-form optimal strategy of the broker. Section 4.2.1 employs the learning equations of Section 4.1.1, i.e., when she employs the midprice as the source of information, and Section 4.2.2 employs the learning equations of Section 4.1.2, i.e., when she employs the trading speed of the informed trader as the source of information. For the first, the filtration of the broker is $({\mathcal{F}}^{B}_{t})_{t\in{\mathfrak{T}}}$, and for the second, the filtration is $(\mathcal{F}^{B,{\textup{alt}}}_{t})_{t\in{\mathfrak{T}}}$.
In the numerical experiments below, we study the performance of the broker’s strategy when she estimates the private signal according to (4.4)–(4.5) and also (4.9)–(4.10).
The uninformed order flow $\xi$ is mean-reverting around zero and we write its dynamics as
@@ 4.2  The control problems of the broker
$\mathrm{d}X^{B,\nu}_{t}$ $=\big(-\nu_{t}\,\left(S_{t}+\mathfrak{a}\,\nu_{t}\right)+\eta^{*}_{t}\,\left(S_{t}+\mathfrak{b}\,\eta^{*}_{t}\right)+\xi_{t}\,\left(S_{t}+\mathfrak{c}\,\xi_{t}\right)\big)\,\mathrm{d}t\,,\quad X^{B,\nu}_{0}=0\,$, (4.15)
@@
For a given strategy $\nu$, she maximises the performance criterion
@@
where $\mathfrak{y}=(s,x^{B},q^{B},\hat{\alpha},\xi,q^{I})$ for when the broker uses prices as the source of information, and $\mathfrak{y}=(s,x^{B},q^{B},q^{I},\gamma,\hat{\alpha},\xi)$ when the broker uses the trading speed of the informed trader as the source of information. Thus, the control problem that the broker solves is
@@
where the set of admissible control ${\mathcal{A}}^{B}$ depends on the broker’s learning strategies. For the control problem when the broker filters prices, she considers the following set of admissible controls
${\mathcal{A}}^{B}:=\left\{\nu=\left(\nu_{t}\right)_{t\in{\mathfrak{T}}}\,:\,\nu\text{ is $({\mathcal{F}}^{B}_{t})_{t\in{\mathfrak{T}}}-$progressively measurable }\text{and}\,\,\mathbb{E}^{B}\left[\int_{0}^{T}(\nu_{s})^{2}\,ds\right]<\infty\right\}\,.$ (4.18)
When the broker filters the informed trader’s flow, she considers the following set of admissible controls
${\mathcal{A}}^{B}:=\left\{\nu=\left(\nu_{t}\right)_{t\in{\mathfrak{T}}}\,:\,\nu\text{ is $(\mathcal{F}^{B,{\textup{alt}}}_{t})_{t\in{\mathfrak{T}}}-$progressively measurable }\text{and}\,\,\mathbb{E}\left[\int_{0}^{T}(\nu_{s})^{2}\,ds\right]<\infty\right\}\,.$ (4.19)
Here we solve the optimal control for both learning equations.
4.2.1 Solving the control problem when the broker filters prices
In B we provide detailed derivations of the equations that characterise the value function $H$, which is given by
$H(t,\mathfrak{y})=x+q^{B}\,s+G_{0}+2\,G_{1}\,\tilde{\mathfrak{y}}^{\top}+\tilde{\mathfrak{y}}\,G_{2}\,\tilde{\mathfrak{y}}^{\top}$ (4.20)
where $\tilde{\mathfrak{y}}=\begin{pmatrix}q^{B}&\hat{\alpha}&\xi&q^{I}\end{pmatrix}$, and $G_{0}:[0,T]\to\mathbb{R}$, $G_{1}:[0,T]\to{\mathcal{M}_{1\times 4}(\mathbb{R})}$, and $G_{2}:[0,T]\to{\mathcal{M}_{4\times 4}(\mathbb{R})}$ are deterministic functions of time reported in the appendix. The notation $\mathcal{M}_{n\times m}(\mathbb{R})$} denotes the space of $n$ by $m$ matrices with values in $\mathbb{R}$. This ansatz helps us characterise the solution to the control problem provided that the Riccati equation that $G_{2}$ satisfies admits a solution. It turns out that there is a region in the parameter space for which this equation admits a unique solution.
The next lemma shows existence and uniqueness of a solution to the differential equation in $G_{2}$. We assume that the values of the model parameters are such that $1+\mathfrak{b}\,f_{3}(t)>0$ for all $t\in[0,T]$. Note that $\mathfrak{b}\,f_{3}/2=g_{2}^{I}$, thus, the inequality can be written as $g_{2}^{I}>-1/2$. Given $g_{2}^{I}$ satisfies the ODE in (A.4), we see that if the value of the temporary impact and terminal inventory risk parameters for the informed trader ($\mathfrak{b}$, $\beta_{0}^{I}$, and $\beta_{1}^{I}$) are small, this condition is met. In the numerical examples below the values of the model parameters are such that this condition is satisfied.
Proposition 2.
Assume that the values of model parameters are such that $1+\mathfrak{b}\,f_{3}(t)>0$ for all $t\in[0,T]$. There exists $\delta>0$ such that for all $\mathfrak{p}\in[0,\delta)$, the matrix Riccati equation (B.67) admits a unique solution $G_{2}:[0,T]\to{\mathcal{M}_{4\times 4}(\mathbb{R})}$ with terminal condition (B.68).
Once the existence and uniqueness of $G_{2}$ is established, the existence and uniqueness of $G_{0}$ and $G_{1}$ is straightforward. The following theorem, whose proof is in B.5, solves the adaptive control problem of the broker.
Theorem 2.
Let $\mathfrak{p}\in[0,\delta)$ with $\delta$ as in Proposition 2. Then, the HJB equation (B.41) admits a unique solution in $C^{1,2}([0,T],\mathbb{R}^{|\mathfrak{y}|})$ with quadratic growth in $\mathfrak{y}$. Moreover, the broker’s Markovian optimal trading rate for the control problem (4.17) in feedback form is given by
where $\tilde{\mathfrak{y}}_{t}=\begin{pmatrix}Q^{B*}_{t}&\hat{\alpha}_{t}&\xi_{t}&Q^{I}_{t}\end{pmatrix}$, and $P_{7},P_{8}\in{\mathcal{M}_{1\times 4}(\mathbb{R})}$ are in (B.61) and (B.63) in the Appendix.
4.2.2 Solving the control problem when the broker filters the trading speed
In B we provide detailed derivations of the equations that characterise the value function $H$, which is given by
$H(t,\tilde{\mathfrak{y}})=x^{B}+q^{B}\,s+h_{0,0}(t)+2\,H_{0,1}(t)\,\tilde{\mathfrak{y}}^{\top}+\tilde{\mathfrak{y}}\,H_{2}(t)\,\tilde{\mathfrak{y}}^{\top}$ (4.22)
where $\tilde{\mathfrak{y}}=\begin{pmatrix}q^{B}&q^{I}&\gamma&\hat{\alpha}&\xi\end{pmatrix}$, and $h_{0,0}:[0,T]\to\mathbb{R}$, $H_{0,1}:[0,T]\to{\mathcal{M}_{1\times 5}(\mathbb{R})}$, and $H_{2}:[0,T]\to{\mathcal{M}_{5\times 5}(\mathbb{R})}$ are deterministic functions of time reported in B.7.
The existence of a classical solution to the HJB equation for this problem depends on a $3\times 3$ matrix Riccati equation reported in (B.105).
Theorem 3.
If there is a solution to the matrix Riccati equation (B.105), the HJB equation (B.77) admits a unique solution in $C^{1,2}([0,T],\mathbb{R}^{|\mathfrak{y}|})$ with quadratic growth in $\mathfrak{y}$. Moreover, the broker’s Markovian optimal trading rate for the control problem (4.17) in feedback form is given by
$\nu^{*}_{t}={\mathfrak{h}}_{0}(t)+(\mathfrak{p}+{\mathfrak{h}}_{1}(t))\,Q^{B}_{t}+{\mathfrak{h}}_{2}(t)\,Q^{I}_{t}+{\mathfrak{h}}_{3}(t)\,\gamma_{t}+{\mathfrak{h}}_{4}(t)\,\hat{\alpha}^{\text{alt}}_{t}+{\mathfrak{h}}_{5}(t)\,\xi_{t}\,$, (4.23)